# DRtool: An Interactive Tool for Analyzing High-Dimensional Clusterings

### v2

@@ Abstract
Technological advances have spurred an increase in data complexity and dimensionality. We are now in an era in which data sets containing thousands of features are commonplace. To digest and analyze such high-dimensional data, dimension reduction techniques have been developed and advanced along with computational power. Of these techniques, nonlinear methods are most commonly employed because of their ability to construct visually interpretable embeddings. Unlike linear methods, these methods non-uniformly stretch and shrink space to create a visual impression of the high-dimensional data. Since capturing high-dimensional structures in a significantly lower number of dimensions requires drastic manipulation of space, nonlinear dimension reduction methods are known to occasionally produce false structures, especially in noisy settings. In an effort to deal with this phenomenon, we developed an interactive tool that enables analysts to better understand and diagnose their dimension reduction results. It uses various analytical plots to provide a multi-faceted perspective on results to determine legitimacy. The tool is available via an R package named DRtool.
Keywords: Non-Linear Dimension Reduction, R Package, R Shiny
@@ 1 Introduction
The potency of nonlinear dimension reduction methods lies in their flexibility, allowing them to model complex data structures. That same flexibility, however, makes them difficult to use and interpret. Each method requires a slew of hyperparameters that need to be calibrated, and even when adequately calibrated, these methods require a trained eye to interpret. For example, the two most popular nonlinear dimension reduction methods, t-SNE and UMAP, sometimes generate misleading results (Coenen and Pearce, 2024; Wattenberg et al., 2016). The results often cluster, even when no clusters exist in the data, and cluster sizes/locations can be unreliable. We have developed an interactive tool that analysts may use to conduct a post-hoc analysis of their high-dimensional clustering. The tool uses the minimum spanning tree (MST) to describe the global structure of clusters and provide an additional perspective on inter-cluster relationships. This allows analysts to extract more information from their dimension reduction results by making it easier to differentiate the signal and the noise.
In this paper, we describe the analytical plots provided by the tool (Section 2). We present a MST stability experiment, demonstrating the MST’s ability to approximate high-dimensional structure, as well as power and size analyses for a novel hypothesis test (Section 3). And we walk through use of the tool on two separate data sets (Section 4).
@@ 2.1 Minimum Spanning Tree
Graphs have been applied to many multivariate statistical problems. The authors of Rozál and Hartigan (1994) introduced the minimal ascending path spanning tree as a way to test for multimodality. The Friedman-Rafsky test (Friedman and Rafsky, 1979), along with its modern variations (Bhattacharya, 2019; Chen and Friedman, 2017; Chen et al., 2018), use the MST to construct a multivariate two-sample test. Single-linkage clustering (Gower and Ross, 1969) and runt pruning (Stuetzle, 2003) are both intimately related to the MST. In the context of dimension reduction, IsoMap (Tenenbaum et al., 2000) makes use of neighborhood graphs, King and Tidor (2009) introduces the maximum information spanning tree, and Probst and Reymond (2020) uses the MST. These methods, which fall under the category of manifold learning, use graphs to model high-dimensional data assumed to be drawn uniformly from a high-dimensional manifold. An accurate low-dimensional embedding can then be constructed from these graphs. It’s apparent that graphs are useful for describing high-dimensional data, especially when it comes to dimension reduction and cluster analysis. Our tool uses the MST to analyze the reliability of visualizations produced by nonlinear dimension reduction methods.
We’ve opted for the MST for a couple of key properties. Firstly, the MST and shortest paths along it are quick to compute. Secondly, the MST contains a unique path between any two vertices, providing a well-defined metric on the data. Lastly, it provides a good summary of the data’s structure. It contains as a subgraph the nearest-neighbor graph, and any edge deletion in the MST partitions the vertices into two sets for which the deleted edge is the shortest distance between them (Friedman and Rafsky, 1979).
@@ 2.1.1 Simplified Medoid Subtree
To derive the inter-cluster relationships from the MST, we first take the medoid subtree, i.e. the minimal subtree containing the medoid of each cluster, then apply a simplification procedure (Algorithm 1). The algorithm collapses paths of non-medoid vertices into single edges of equal length. We refer to the output as the simplified medoid subtree. It encodes the global inter-cluster relationships within the data.
@@
MST T=(V,E)T=(V,E)with cluster medoids $m_{1},\ldots,m_{k}\in V$
[…]
2:Let $v\in V^{\prime}\setminus\{m_{1},\ldots,m_{k}\}$ with $deg(v)=2$ and neighbors $a,b\in V^{\prime}$. Let $d(v,a)$ and $d(v,b)$ be the weights of the edges incident to $v$.
[…]
@@ 2.1.2 Robinson-Foulds Metric
To compare simplified medoid subtrees, we use the Robinson-Foulds metric (Robinson and Foulds, 1981). The R-F metric was originally introduced to quantify the dissimilarity of phylogenetic trees, but the algorithm generalizes to arbitrary weighted trees. It looks at partitions of each tree created by removing individual edges, then counts the number of partitions present in one tree but not the other. We modified the algorithm (Algorithm 2) to specifically measure the dissimilarity in medoid vertices.
@@
[…]
@@
[…]
$G\Leftarrow(V_{2},E_{2}\setminus\{e\})$ with connected components $G_{1}$ and $G_{2}$
[…]
$M_{2}\Leftarrow\{n_{1},\ldots,n_{k}\}\cap V(G_{2})$
[…]
@@ 2.2 The Tool
The main objective is to analyze and leverage the structural data embedded in the MST. For example, paths between clusters are used to study inter-cluster relationships in the context of the underlying manifold from which the data are drawn.
@@ 2.3.1 The PCA – rCCA Method
Let $P\in\mathbb{R}^{k\times p}$ be the matrix of high-dimensional path points with endpoints $p_{1},p_{k}\in\mathbb{R}^{p}$. Let $X\in\mathbb{R}^{n\times p}$ be the matrix of points of interest, i.e. the points belonging to the selected groups and the points along the path.
@@
The first two canonical variate pairs are used to construct a two-dimensional projection that maximizes the covariance between the projections of $P$ and $P_{d}$. This process generates a two-dimensional subspace on which we can project all of $X$. Regularization is required to avoid singularity because $p$ is often much greater than $k$. The regularization constant for $P$ is chosen using cross-validation. No regularization constant is needed for $P_{d}$. See Tuzhilina et al. (2023) for details.
One issue with this method is the projected path often travels along the outskirts of the plot. This is due to the near-orthogonality of high-dimensional data (Diaconis and Freedman, 1984). Because the non-path points are often nearly orthogonal with the projection subspace, they are overly shrunk in the projection. The path points are less affected because the projection subspace is selected to retain the path’s shape. While this phenomenon doesn’t discredit the entire plot, it leads to misrepresentation of the path’s location relative to the rest of the points.
To alleviate this issue, we apply PCA on the entirety of $X$ to prior to applying rCCA. Removing extraneous dimensions containing mostly noise limits the confusion of excess noise for independence. When rCCA is applied post-PCA, the projected path’s relative position to the rest of the points is more credible.
@@ 2.3.2 Calibrating Hyperparameters
The user is responsible for calibrating the dimensionality of the PCA step and the degree $d$ of the reference polynomial design matrix. To pick a number of dimensions, the user is recommended to start with a moderately large number, relative to the dimensionality of the original data. The proportion of variance retained in the selected number of dimensions is conveniently displayed in the upper righthand corner of the plot. A larger number of dimensions retains more information, but may misrepresent the location of the path relative to the rest of the points, while a smaller number of dimensions may diminish some of the variation in the data. As such, the user is encouraged to try different numbers of dimensions. To calibrate $d$, it is recommended to start with $d=2$ then increment $d$ until the shape of the path stabilizes.
The user is also given the option to overlay a kernel density estimate. In order to do so, the bandwidth must be calibrated. The recommend procedure is to begin with a large bandwidth that estimates one mode, then gradually decrease the bandwidth until two modes appear. If the two modes correspond with the two groups of interest, and more modes do not immediately appear when continuing to decrease the bandwidth, then a bimodal distribution is a reasonable way to describe the data.
@@ 2.4.1 The Test Statistic
The test statistic, meant to quantify local connectivity, is based on the number of edges connecting the two groups of interest. However, counting single edges is too restrictive of a measure. Consider the case when the two selected groups are polar ends of the same cluster. Because the medial region of the cluster does not belong to either group, there will be zero edges connecting the two groups, indicating the maximal degree of separation. This result is undesirable because the two groups actually belong to the same cluster.
Instead, the test statistic counts the number of connecting paths rather than single edges. These paths are referred to as crossings and are counted according to the following procedure. The minimal subtree containing both groups is isolated. Because the two groups may not be adjacent in the MST, this subtree may contain points belonging to other clusters as well. To extract the structural relationship between the two groups of interest, the subtree must be simplified. The simplification process collapses paths between the two groups of interest into edges that can be counted (Algorithm 3).
@@
[…]
@@
[…]
@@
[…]
To count the number of crossings, the number of edges between the two groups in the simplified subtree are counted. It is also possible for a point of non-interest to act as a mediator along a path between the two groups of interest. To account for this scenario, for each point of non-interest adjacent to both groups, we also count its maximal degree to both groups.
@@ 2.4.2 The Null Distribution
The null distribution should correspond to the number of crossings in the case when both groups belong to the same cluster. Because a cluster can be drawn from a number of unimodal distributions, we must consider a composite null hypothesis including all such distributions. Among these distributions, the distribution that maximizes the probability of rejection must be used to ensure the test has the correct size. That way, the probability of Type I error does not exceed the pre-specified significance level under any other member of the composite null hypothesis as well.
@@
for $j=1,2$. The product of singular values is used to estimate the volume of each group because high-dimensional clusters tend to look Gaussian (Diaconis and Freedman, 1984). Extraneous noise dimensions are removed prior to this process to avoid biasing the volume estimates. Now suppose $D_{1}<D_{2}$. Then $n_{1}$ points are uniformly sampled from a hyperrectangle with side lengths $\sqrt{12}\sigma_{i}^{1}$ (the factor $\sqrt{12}$ ensures the variance of the sample in each principle direction is equal to the variance of the original data in each principal component), then the number of edges crossing the hyperplane in the MST is recorded. Repeated simulation yields an approximate null distribution to which the test statistic is compared. The returned $p$-value is the percentile of the test statistic within this bootstrapped null distribution. A one-sided test is employed because we are only interested in rejecting the null for sufficiently small numbers of edge crossings.
Power and size analyses are conducted in controlled situations. They are described in Section 2.7, and the results are presented in Section 3.2.
@@ 3.1 MST Stability Experiment
To demonstrates the MST’s robust ability to describe global structure, we conducted a stability experiment. 1,500 samples were randomly chosen from the MNIST data set of handwritten digits (Deng, 2012). Each $28\times 28$-pixel image was flattened into a vector of length $28^{2}=784$, so the data contain 1,500 samples in $784$ dimensions. A PCA pre-processing step was employed to reduce the number of dimensions to 300. The simplified medoid subtree $T$ was then calculated.

### v3

@@ Abstract
When faced with new data, we often conduct a cluster analysis to obtain a better understanding of the data’s structure and the archetypical samples present in the data. This process often includes visualization of the data, either as a way to discover or verify clusters. However, the increases in data complexity and dimensionality has made this step very tricky. To visualize data, nonlinear dimension reduction methods are the de facto standard for their ability to non-uniformly stretch and shrink space in order to preserve local clusters. Because this process requires a drastic manipulation of space, however, nonlinear dimension reduction methods are known to produce false structures, especially when mishandled. A common consequence that often goes undetected by the untrained eye is over-clustering of the data. In efforts to deal with this phenomenon, we developed an interactive tool that empowers analysts to distinguish false clusters and better interpret their high-dimensional clustering results. The tool uses various analytical plots to provide a multi-faceted perspective on the data’s global structure as well as local inter-cluster relationships, helping users determine the legitimacy of their high-dimensional clustering results. The tool is available via an R package named DRtool.
Keywords: Cluster Analysis, Non-Linear Dimension Reduction, R Package
@@ 1 Introduction
The potency of nonlinear dimension reduction methods lies in their flexibility, allowing them to model complex data structures. That same flexibility, however, makes them difficult to use and interpret. Each method requires a slew of hyperparameters that need to be calibrated, and even when adequately calibrated, these methods require a trained eye to interpret. For example, the two most popular nonlinear dimension reduction methods, t-SNE and UMAP, sometimes generate misleading results (understanding_UMAP; Distill). The results often cluster, even when no clusters exist in the data, and cluster sizes/locations can be unreliable. This is especially problematic in the context of high-dimensional cluster analysis when the tools for determining cluster legitimacy are limited. To add to the toolbox, we have developed an interactive tool that analysts may use to conduct a post-hoc analysis of their high-dimensional clustering. The tool uses the minimum spanning tree (MST) to provide a broad view of the data’s global structure as well as local perspectives on inter-cluster relationships via various plots and tests. This allows analysts to extract more information from their dimension reduction results by making it easier to differentiate the signal from the noise.
In this paper, we describe the analytical plots provided by the tool (Section 3). We present a MST stability experiment, demonstrating the MST’s ability to approximate high-dimensional structure, as well as power and size analyses for a novel hypothesis test (Section 4). And we walk through the use of the tool on two separate data sets (Section 5).
2 Related Works
Past works have investigated both the difficulty in interpreting non-linear dimension reduction results and the consequences specific to cluster analysis, but there does not yet exist an interactive tool focused specifically on high-dimensional clustering. The extent of interactive tools is the grand tour method (grand_tour), along with its modern variations (tour1; tour2), which provide an interactive sequence of images describing high-dimensional data from various perspectives. These tools, however, do not focus on micro inter-cluster relationships. Other works address the difficulties of high-dimensional clustering (clust1; clust2), but they are singular methods, as opposed to interactive toolboxes. Our tool fills the void of a comprehensive tool specific to diagnosing high-dimensional clusterings that is intuitive, interactive, and accessible.
The use of graphs, and specifically the MST, to model high-dimensional data has also been well-documented and is explained in Section 3.1.
@@ 3.1 Minimum Spanning Tree
Graphs have been applied to many multivariate statistical problems. The authors of MAP_test introduced the minimal ascending path spanning tree as a way to test for multimodality. The Friedman-Rafsky test (Friedman-Rafsky-test), along with its modern variations (Friedman-Rafsky1; Friedman-Rafsky2; Friedman-Rafsky3), use the MST to construct a multivariate two-sample test. Single-linkage clustering (single-linkage) and runt pruning (runt_pruning) are both intimately related to the MST. In the context of dimension reduction, IsoMap (IsoMap) makes use of neighborhood graphs, MIST_example introduces the maximum information spanning tree, and MST_example uses the MST. These methods, which fall under the category of manifold learning, use graphs to model high-dimensional data assumed to be drawn uniformly from a high-dimensional manifold. An accurate low-dimensional embedding can then be constructed from these graphs. It’s apparent that graphs are useful for describing high-dimensional data, especially when it comes to dimension reduction and cluster analysis. Our tool uses the MST to analyze the reliability of visualizations produced by nonlinear dimension reduction methods.
We’ve opted for the MST for a couple of key properties. Firstly, the MST and shortest paths along it are quick to compute. Secondly, the MST contains a unique path between any two vertices, providing a well-defined metric on the data. Lastly, it provides a good summary of the data’s structure. It contains as a subgraph the nearest-neighbor graph, and any edge deletion in the MST partitions the vertices into two sets for which the deleted edge is the shortest distance between them (Friedman-Rafsky-test).
@@ 3.1.1 Simplified Medoid Subtree
To encode the inter-cluster relationships present in the MST, we define the simplified medoid subtree as follows. We take the medoid subtree, i.e. the minimal subtree containing the medoid of each cluster, then apply a simplification procedure (Algorithm 1). The algorithm collapses paths of non-medoid vertices into single edges of equal length by iteratively replacing degree-2 vertices with a single edge. We refer to the output as the simplified medoid subtree. It encodes the global inter-cluster relationships within the data. See Figure 1 for an example.
@@
1: MST $T=(V,E)$ with vertex set $V$, edge set $E$, and cluster medoids $m_{1},\ldots,m_{k}\in V$
2: $T^{\prime}=(V^{\prime},E^{\prime})\Leftarrow$ minimal subtree of $T$ containing all $m_{i}$
3: repeat
4:   Let $v\in V^{\prime}\setminus\{m_{1},\ldots,m_{k}\}$ with degree $deg(v)=2$ and neighbors $a,b\in V^{\prime}$. Let $d(v,a)$ and $d(v,b)$ be the weights of the edges incident to $v$.
5:   Replace $v$ and its two incident edges with an edge connecting $a$ and $b$ with weight $d(v,a)+d(v,b)$.
6: until $T^{\prime}$ no longer contains non-medoid vertices with degree two.
7: output T’
Figure 1: Example of Algorithm 1. The first step is to compute the medoid subtree $T^{\prime}$. Then non-medoid degree-2 vertices are replaced with single edges.
@@ 3.1.2 Robinson-Foulds Metric
To compare simplified medoid subtrees, we use the Robinson-Foulds metric (RF_metric). The R-F metric was originally introduced to quantify the dissimilarity of phylogenetic trees, but the algorithm generalizes to arbitrary weighted trees. It looks at partitions of each tree created by removing individual edges then counts the number of partitions present in one tree but not the other. We modified the R-F metric to specifically measure the dissimilarity in medoid vertices. In particular, Algorithm 2 compares the partitions of the medoid vertices created by removing individual edges, as opposed to partitions of the entire vertex set.
@@
2: $P_{1}\Leftarrow\{\}$
for $e\in E_{1}$ do
4:   $G\Leftarrow(V_{1},E_{1}\setminus\{e\})$ with connected components $G_{1}$ and $G_{2}$
@@
6:   $M_{2}\Leftarrow\{m_{1},\ldots,m_{k}\}\cap V(G_{2})$
   $P_{1}\Leftarrow\textsc{Add}(P_{1},\{M_{1},M_{2}\})$
8: $P_{2}\Leftarrow\{\}$
for $e\in E_{2}$ do
10:   $G\Leftarrow(V_{2},E_{2}\setminus\{e\})$ with connected components $G_{1}$ and $G_{2}$
   $M_{1}\Leftarrow\{n_{1},\ldots,n_{k}\}\cap V(G_{1})$
12:   $M_{2}\Leftarrow\{n_{1},\ldots,n_{k}\}\cap V(G_{2})$
   $P_{2}\Leftarrow\textsc{Add}(P_{2},\{M_{1},M_{2}\})$
14: output $\frac{\left|P_{1}\Delta P_{2}\right|}{2\left|P_{1}\cap P_{2}\right|}$
@@ 3.2 The Tool
The main objective is to analyze and leverage the structural data embedded in the MST. For example, paths between clusters are used to study inter-cluster relationships in the context of the underlying manifold from which the data are assumed to be drawn.
@@
From now on, “group” will refer to the user-selected groups, while “cluster” will refer to the clusters in the user-inputted clustering.
@@ 3.3.1 The PCA – RCCA Method
Let $p_{1},\ldots,p_{k}\in\mathbb{R}^{p}$ be the points along the path and
$P=\begin{bmatrix}\text{---}&p_{1}&\text{---}\\ &\vdots&\\ \text{---}&p_{k}&\text{---}\end{bmatrix}\in\mathbb{R}^{p\times k}$
the matrix of path points.
@@
In doing so, we are linearly transforming the path so that its shape best matches that of a degree $d$ parametric polynomial. The first two canonical variate pairs are used to construct a two-dimensional projection that maximizes the covariance between the projections of $P$ and $P_{d}$. This process generates a two-dimensional projection subspace onto which we can project the path and both groups.
In most cases, however, regularization is required to avoid singularity because $p$ is often much greater than $k$. RCCA (RCCA) adds a regularization constant along the diagonals of each covariance matrix to ensure nonsingularity. The regularization constant for $P$ is chosen via cross-validation, and no regularization constant is needed for $P_{d}$.
One issue with this method is the projected path often travels along the outskirts of the plot. This is due to the near-orthogonality of high-dimensional data (near-orthogonal). Because the non-path points are often nearly orthogonal with the projection subspace, they are overly shrunk in the projection. The path points are less affected because the projection subspace is selected to retain the path’s shape. While this phenomenon doesn’t discredit the entire plot, it leads to misrepresentation of the path’s location relative to the rest of the points.
To alleviate this issue, we apply PCA on the entirety of $X$ prior to applying RCCA. Removing extraneous dimensions containing mostly noise limits the confusion of excess noise for independence. When RCCA is applied post-PCA, the projected path’s relative position to the rest of the points is more credible.
@@ 3.3.2 Calibrating Hyperparameters
The user is responsible for calibrating the dimensionality of the PCA step and the degree $d$ of the reference polynomial design matrix. To pick a number of dimensions, the user is recommended to start with a moderately large number, relative to the dimensionality of the original data. The proportion of variance retained in the selected number of dimensions is conveniently displayed in the upper righthand corner of the plot. A larger number of dimensions retains more information but may misrepresent the location of the path relative to the rest of the points, while a smaller number of dimensions may diminish some of the variation in the data. As such, the user is encouraged to try different numbers of dimensions. To calibrate $d$, it is recommended to start with $d=2$ then increment $d$ until the shape of the path stabilizes.
The user is also given the option to overlay a kernel density estimate. In order to do so, the bandwidth must be calibrated. The recommended procedure is to begin with a large bandwidth that estimates one mode, then gradually decrease the bandwidth until two modes appear. If the two modes correspond with the two groups of interest, and more modes do not immediately appear when continuing to decrease the bandwidth, then a bimodal distribution is a reasonable way to describe the data.
@@ 3.4.1 The Test Statistic
The test statistic, meant to quantify local connectivity, is based on the number of edges connecting the two groups of interest. However, counting single edges is too restrictive of a measure. Consider the case when the two selected groups are polar ends of the same cluster. Because the medial region of the cluster does not belong to either group, there will be zero edges connecting the two groups, indicating the maximal degree of separation. This result is undesirable because the two groups actually belong to the same cluster in this hypothetical scenario.
Instead, the test statistic counts the number of connecting paths rather than single edges. These paths are referred to as crossings and are counted according to the following procedure. The minimal subtree containing both groups is isolated. Because the two groups may not be adjacent in the MST, this subtree may contain points belonging to other clusters as well. To extract the structural relationship between the two groups of interest, the subtree must be simplified. The simplification process collapses paths between the two groups of interest into edges that can be counted (Algorithm 3). See Figure 2 for an example.
@@
$T^{\prime}=(V^{\prime},E^{\prime})\Leftarrow$ minimal subtree of $T$ containing $V_{1}\cup V_{2}$
3: repeat
@@
   Replace $v$ and its two incident edges with an edge connecting $a$ and $b$ with weight $d(v,a)+d(v,b)$.
6: until $T^{\prime}$ no longer contains non-group vertices with degree two.
repeat
@@
9:   Collapse the edge connecting $v_{1}$ and $v_{2}$. The combined vertex is adjacent to all neighbors of $v_{1}$ and $v_{2}$.
until $T^{\prime}$ no longer adjacent non-group vertices.
Figure 2: Example of Algorithm 3. First, degree-2 vertices belonging to neither group are replaced with single edges. Then adjacent vertices belonging to neither group are collapsed into a single vertex.
To count the number of crossings, the number of edges between the two groups in the simplified subtree is counted. It is also possible for a point of non-interest to act as a mediator along a path between the two groups of interest (see the unlabeled vertex in the third panel of Figure 2). To account for this scenario, for each point of non-interest adjacent to both groups, we also count its maximal degree to both groups.
@@ 3.4.2 The Null Distribution
The null distribution should correspond to the number of crossings in the case when both groups belong to the same cluster. Because a cluster can be drawn from a number of unimodal distributions, we must consider a composite null hypothesis including all such distributions. Among these distributions, the distribution that maximizes the probability of rejection must be used to ensure the test has the correct size. That way, the probability of Type I error does not exceed the pre-specified significance level under any member of the composite null hypothesis.
@@
for groups $j=1,2$. The product of singular values $\sigma_{i}^{j}$ is used to estimate the volume of each group because high-dimensional clusters tend to look Gaussian (near-orthogonal). Extraneous noise dimensions are removed prior to this process to avoid biasing the volume estimates. Now suppose $D_{1}<D_{2}$. Then $n_{1}$ points are uniformly sampled from the hyperrectangle
$R=\prod_{i=1}^{p}\left[-\frac{\sqrt{12}}{2}\sigma_{i}^{1},\frac{\sqrt{12}}{2}\sigma_{i}^{1}\right].$
Note, $R$ has side lengths $\sqrt{12}\sigma_{i}^{1}$ so that the marginal distribution along the $i^{\textrm{th}}$ dimension has standard deviation $\sigma_{i}^{1}$, i.e. the standard deviation of the $i^{\textrm{th}}$ principal component of the original data. From there, the number of edges in the MST crossing the $x_{1}=0$ hyperplane is recorded. Repeated simulation of this procedure yields an approximate null distribution to which the test statistic is compared. The returned $p$-value is the percentile of the test statistic within this bootstrapped null distribution. A one-sided test is employed because we are only interested in rejecting the null for sufficiently small numbers of edge crossings.
See Section 4.2 for power and size analyses.
@@ 4.1 MST Stability Experiment
To demonstrates the MST’s robust ability to describe global structure, we conducted a stability experiment. 1,500 samples were randomly chosen from the MNIST data set of handwritten digits (MNIST). Each $28\times 28$-pixel image was flattened into a vector of length $28^{2}=784$, so the data contain 1,500 samples in $784$ dimensions. A PCA pre-processing step was employed to reduce the number of dimensions to 300. The simplified medoid subtree $T$ was then calculated.